# Combining Neural Networks and Symbolic Regression for Analytical Lyapunov Function Discovery

### v1

@@ Abstract
We propose CoNSAL (Combining Neural networks and Symbolic regression for Analytical Lyapunov function) to construct analytical Lyapunov functions for nonlinear dynamic systems. This framework contains a neural Lyapunov function and a symbolic regression component, where symbolic regression is applied to distill the neural network to precise analytical forms. Our approach utilizes symbolic regression not only as a tool for translation but also as a means to uncover counterexamples. This procedure terminates when no counterexamples are found in the analytical formulation. Compared with previous results, our algorithm directly produces an analytical form of the Lyapunov function with improved interpretability in both the learning process and the final results. We apply our algorithm to 2-D inverted pendulum, path following, Van Der Pol Oscillator, 3-D trig dynamics, 4-D rotating wheel pendulum, 6-D 3-bus power system, and demonstrate that our algorithm successfully finds their valid Lyapunov functions.
@@ 2.1 Learning-based Lyapunov function construction
The area of learning-based Lyapunov function construction is experiencing significant growth. A prominent model in this field is the Neural Lyapunov Control (Chang et al., 2019). This framework includes a neural-network-based Lyapunov function and a learnable linear controller, with stability verified by an SMT solver. This methodology was extended by (Zhou et al., 2022), which introduced a neural network to model unknown dynamics and another for control. Additionally, (Dai et al., 2021; Wu et al., 2023) consider discrete-time systems and neural network controllers, verifying stability via MIP solvers. One limitation of these methods is the need for linearization of dynamics or other relaxations for formal verification. (Yang et al., 2024) deploys $\alpha,\beta$-CROWN for neural network verification due to its scalability, and extends the previous control scheme from state feedback to output feedback control. However, these formal verification methods face challenges with large neural networks: SMT solvers can handle up to 30 neurons, MIP solvers are limited to networks with up to 200 neurons (Dawson et al., 2023), and (Yang et al., 2024) has 16 neurons each layer for the Lyapunov neural network.
@@ 4.2 Symbolic Regression and Falsification
Without a perfect root-finder, our method accelerates counterexample generation, leading to an analytical Lyapunov function. Formal validation of the final function may still be required even if numerical methods fail to find roots, which we defer to future work.
@@ 4.3 Compositional Neural Lyapunov Function
To address this issue, an intuitive way is to find compositional Lyapunov functions for the high-dimensional system instead of a single Lyapunov function, i.e. finding one Lyapunov function $V_{i}(x_{i})$ for each subsystem $i$. Moreover, if the networked system has a similar symmetric structure for different subsystems, the local Lyapunov function can share the same structure across different subsystems. Following (Zhang et al., 2023), when the individual Lyapunov functions $V_{i}(x_{i})$ satisfy an Input-to-State Stability (ISS)-style condition, they will certify the stability of the entire dynamical system.
@@
Figure 3: Lie Derivative of the Neural Lyapunov Function under linear path following dynamics, updated with the proposed algorithm. Showing the result of the first zero epochs.
@@ 5 Experiment
In this section, we validate the proposed algorithm with various nonlinear dynamics by finding their Lyapunov functions, where the autonomous systems (without additional control) are stable at the origin. We test the proposed algorithm using the neural Lyapunov network (5) on 2-D path following, inverted pendulum, Van Der Pol oscillator, 3-D trig dynamics, 4-D rotating wheel pendulum, and 6-D nonlinear dynamics. The compositional neural Lyapunov function is applied to a 3-bus power system. To ease the requirement for symbolic regression to find the exact constants in the Lyapunov functions and accommodate the error of numerical root-finding tools, we allow a violation of the stability conditions less than $10^{-4}$ in value and the Lyapunov function is only verified at a given local region numerically. This approach aims to find an analytically correct Lyapunov function with reasonable constants’ errors. Global stability can be assessed through further inspection. In this work, multiple initial guesses are applied to the root finder to facilitate root finding.
We first demonstrate the evolution of the Neural Lyapunov function, and compare how the proposed framework optimizes the neural Lyapunov function with the SMT-based scheme (Chang et al., 2019) in the 2D path following environment, where our approach can achieve better performance with respect to the SMT solver. The runtime information and the found analytical Lyapunov functions are summarized in Table 2. Details about dynamics are available in Appendix A. Neural network configurations and base operator selection are available in Appendix B.
@@ 5.1 Evolution of Neural Lyapunov Function
Linear path following is a classic nonlinear control task, where a wheeled robot is controlled to follow a linear path with dynamics (7) given by (Samson, 1992):
@@
We further compare our framework with the SMT-based neural Lyapunov function (Chang et al., 2019), where the SMT solver is deployed to generate counterexamples for the neural network and validate the stability conditions. We use the same two-layer neural networks with 128 hidden units for both methods. Given the relatively large neural network, the SMT-based approach can only achieve 30% successful rate for converging, while ours achieved 80 % successful rate for finding the valid symbolic Lyapunov function. We compare their performance when they successfully converge. Following Table 1, due to the complexity of the neural network, the SMT solver consumes much more time compared with symbolic-regression-based verification.
@@
The counterexamples generated by both algorithms are presented in Figure 4. Following this result, the SMT solver can identify a small region such that samples from the region will violate the stability constraint. Thus counterexamples are generated by randomly sampling in this region. However, this method provides counterexamples focusing on this small region and can potentially lead to overfitting. Recall that with PySR, multiple symbolic expressions are generated at each complexity level. The roots of these formulations can be efficiently found numerically, which are illustrated in the middle plot of Figure 4. With acceptable approximating errors, the roots outline the region of violation and the counterexamples can be efficiently generated near the roots. In this case, our approach finds counterexamples covering most of the violation areas in the state space and facilitates the successful discovery of a valid Lyapunov function.
@@
  • Note: Verification time represents the average time consumption for each single-step verification. Only success trials’ convergence time and convergence epoch are included in this table. The success rate is calculated out of 10 different trials. Experiment which found valid Lyapunov function within 100 epochs is considered as a successful trail.
@@
Dynamics Time Epoch Found Lyapunov Functions Stability Succ %
@@
*. $E$ is the energy function of the corresponding dynamics, which is defined at A.4.
$\dagger$. In this column, ‘g.a.s’ represents globally asymptotically stable, ‘l.a.s.’ represents locally asymptotically stable.
@@ 5.2 Runtime and Found Analytical Lyapunov Functions
We summarize the runtime and the found Lyapunov functions in table 2. In general, the proposed framework can achieve more than $60\%$ success rate for all the tested dynamics and find at least a local Lyapunov function within $50$ symbolic regression calls, where a local Lyapunov function means the Lyapunov function is valid in the tested local region. Global stability is further extended by human experts. Notably, our algorithm demonstrates fast convergence for the tested 4-D and 6-D systems. We also find that our method can occasionally discover new Lyapunov functions valid locally (detailed in A.2). Here, ‘new’ means that the Lyapunov function is not included in the existing literature.
However, it gets the lowest success rate with the 2-D inverted pendulum system and the 6-D 3-bus power system. The challenge comes from optimizing the constants in the Lyapunov function based on given physical constants in the dynamics. For instance, consider the inverted pendulum system, its local Lyapunov function can be represented as $V(x_{1},x_{2})=(1-cos(x_{1}))+\frac{l}{2g}x_{2}^{2}$, where $x_{1}$ is the angular position from the inverted position, $x_{2}$ is the angular velocity, and $g,m,l,b$ are acceleration of gravity, mass of inverted object, the length of string, and the coefficient of friction respectively. Setting $l=5$\mathrm{m}$$, $g=9.81m/s^{2}$, it is challenging for the symbolic regression to find the correct coefficients. In failed cases, our framework often finds the correct structure but incorrect coefficients, e.g. $V=1-cos(x_{1})+0.26x_{2}^{2}$, which is not valid.
In cases where finding a Lyapunov function is challenging but prior system knowledge is available, we incorporate this knowledge into the Lyapunov function design. Specifically, for the 4-D rotating wheel pendulum, we use the energy function $E$ as part of the input to the neural Lyapunov function, along with all state variables. As a result, a Lyapunov based on the energy function is revealed by our method. Details of all tested dynamics and considered local regions are provided in Appendix A.
@@ 6 Conclusion
In this paper, we propose CoNSAL, a framework to construct analytical Lyapunov functions for stable nonlinear dynamics. Our approach integrates a neural network learner with a symbolic regression solver, which extracts an analytical formula from the neural Lyapunov function. This symbolic Lyapunov function is further validated with numerical root-finding approaches. When the Lyapunov function is invalid, we efficiently sample counterexamples around the roots to facilitate the training process. The resulting valid Lyapunov function can potentially demonstrate global stability or aid experts in designing controllers. Compared to existing work, CoNSAL provides an analytical Lyapunov function instead of neural ones, which significantly improves the interpretability and generalizability of the final results. Moreover, it bypasses the need for neural network verification, enabling the use of larger neural networks and allowing for scalability to more complex systems. The efficiency of the proposed framework is validated in multiple nonlinear dynamical systems ranging from 2-D to 6-D, where our method successfully finds a valid local Lyapunov function for all tested cases.
@@
Appendix A Dynamics of the Tested Systems
A.1 Inverted Pendulum
@@ A.2 Van Der Pol Oscillator
A.4 Rotating Wheel Pendulum
@@ A.6 Power System
Appendix B Neural Network Structure and Symbolic Regression Setting

### v3

@@ Abstract
We propose CoNSAL (Combining Neural networks and Symbolic regression for Analytical Lyapunov function) to construct analytical Lyapunov functions for nonlinear dynamic systems. This framework contains a neural Lyapunov function and a symbolic regression component, where symbolic regression is applied to distill the neural network to precise analytical forms. Our approach utilizes symbolic regression not only as a tool for translation but also as a means to uncover counterexamples. This procedure terminates when no counterexamples are found in the analytical formulation. Compared with previous results, CoNSAL directly produces an analytical form of the Lyapunov function with improved interpretability in both the learning process and the final results. We apply CoNSAL to 2-D inverted pendulum, path following, Van Der Pol Oscillator, 3-D trig dynamics, 4-D rotating wheel pendulum, 6-D 3-bus power system, and demonstrate that our algorithm successfully finds their valid Lyapunov functions. Code examples are available at github.
@@ 2.1 Learning-based Lyapunov function construction
The area of learning-based Lyapunov function construction is experiencing significant growth. A prominent model in this field is the Neural Lyapunov Control (Chang et al., 2019). This framework includes a neural-network-based Lyapunov function and a learnable linear controller, with stability verified by an SMT solver. This methodology was extended by (Zhou et al., 2022), which introduced a neural network to model unknown dynamics and another for control. For switched affine systems, (Feng et al., 2024) introduces a joint learning scheme for a common Lyapunov function and the controller. Additionally, (Dai et al., 2021; Wu et al., 2023) consider discrete-time systems and neural network controllers, verifying stability via MIP solvers. One limitation of these methods is the need for linearization of dynamics or other relaxations for formal verification. (Yang et al., 2024) deploys $\alpha,\beta$-CROWN for neural network verification due to its scalability, and extends the previous control scheme from state feedback to output feedback control. However, these formal verification methods face challenges with large neural networks: SMT solvers can handle up to 30 neurons, MIP solvers are limited to networks with up to 200 neurons (Dawson et al., 2023), and (Yang et al., 2024) has 16 neurons each layer for the Lyapunov neural network.
@@ 4.2 Symbolic Regression and Falsification
Without a perfect root finder, our method accelerates counterexample generation, leading to an analytical Lyapunov function. In experiments, we apply SMT to the found Lyapunov functions for final validation.
@@ 4.3 Compositional Neural Lyapunov Function
To address this issue, an intuitive way is to find compositional Lyapunov functions for the high-dimensional system instead of a single Lyapunov function, i.e. finding one Lyapunov function $V_{i}(x_{i})$ for each subsystem $i$. Moreover, if the networked system has a similar symmetric structure for different subsystems, the local Lyapunov functions can share the same structure across different subsystems. Following (Zhang et al., 2023), when the individual Lyapunov functions $V_{i}(x_{i})$ satisfy an Input-to-State Stability (ISS)-style condition, they will certify the stability of the entire dynamical system.
@@
Figure 3: Lie Derivative of the Neural Lyapunov Function under linear path following dynamics, updated with the proposed algorithm. Showing the result of the first two epochs.
@@ 5 Experiment
In this section, we validate the proposed algorithm with various nonlinear dynamics by finding their Lyapunov functions, where the autonomous systems (without additional control) are stable at the origin. We test the proposed algorithm using the neural Lyapunov network (5) on 2-D path following, inverted pendulum, Van Der Pol oscillator, 3-D trig dynamics, 4-D rotating wheel pendulum, and 6-D nonlinear dynamics. The compositional neural Lyapunov function is applied to a 3-bus power system. To ease the requirement for symbolic regression to find the exact constants in the Lyapunov functions and accommodate the error of numerical root-finding tools, we allow a violation of the stability conditions less than $10^{-4}$ in value and the Lyapunov function is only verified at a given local region numerically. CoNSAL aims to find an analytically correct Lyapunov function with tolerable errors. Global stability can be assessed through further inspection. In this work, multiple initial guesses are applied to the root finder to facilitate root finding.
We first demonstrate the evolution of the Neural Lyapunov function, and compare how the proposed framework optimizes the neural Lyapunov function with the SMT-based scheme (Chang et al., 2019) in the 2D path following environment, where our approach can achieve better performance with respect to the SMT solver. The runtime information and the found analytical Lyapunov functions are summarized in Table 2. Details about dynamics are available in Appendix B. Neural network configurations and base operator selection are available in Appendix C.
@@ 5.1 Evolution of Neural Lyapunov Function
Linear path following is a classic nonlinear control task, which controls a wheeled robot to follow a linear path with dynamics (7) given by (Samson, 1992):
@@
We further compare our framework with the SMT-based neural Lyapunov function (Chang et al., 2019), where the SMT solver is deployed to generate counterexamples for the neural network and validate the stability conditions. We use the same two-layer neural networks with 128 hidden units for both methods. Given the relatively large neural network, the SMT-based approach can only achieve 30% successful rate for converging, while ours achieved 80 % successful rate for finding the valid symbolic Lyapunov function. Following Table 1, due to the complexity of the neural network, the SMT solver consumes much more time compared with symbolic-regression-based verification.
@@
The counterexamples generated by both algorithms are presented in Figure 4. Following this result, the SMT solver can identify a small region such that samples from the region will violate the stability constraint. Thus counterexamples are generated by randomly sampling in this region. However, this method provides counterexamples focusing on this small region and can potentially lead to overfitting. With PySR, multiple symbolic expressions are generated at different complexity levels. The roots of these formulations can be efficiently found numerically, which are illustrated in the middle plot of Figure 4. With acceptable approximating errors, the roots outline the region of violation and the counterexamples can be efficiently generated near the roots. In this case, our approach finds counterexamples covering most of the violation areas in the state space and facilitates the successful discovery of a valid Lyapunov function.
@@
  • Note: Verification time measures the average time consumption for each single-step verification. Only success trials’ convergence time and convergence epoch are included in this table. The success rate is calculated out of 10 different trials. An experiment that found a valid Lyapunov function within 100 epochs is considered a successful trial.
@@
Dynamics Time Epoch Found Lyapunov Functionsa Stability Succ %
@@
$a$. All Found Lyapunov functions passed SMT solver’s verification for Lyapunov conditions, with tolerable error $\epsilon=10^{-4}$.
*. $E$ is the energy function of the corresponding dynamics, which is defined at B.4.
$\ddagger$. In this column, ‘g.a.s’ represents globally asymptotically stable, and ‘l.a.s.’ represents locally asymptotically stable.
@@ 5.2 Runtime and Found Analytical Lyapunov Functions
We summarize the runtime and the found Lyapunov functions in table 2. In general, the proposed framework can achieve more than $60\%$ success rate for all the tested dynamics and find at least a local Lyapunov function within $50$ symbolic regression calls, where a local Lyapunov function means the Lyapunov function is valid in the tested local region. Global stability is further extended by human experts. Notably, our algorithm demonstrates fast convergence for the tested 4-D and 6-D systems. We also find that our method can occasionally discover new Lyapunov functions valid locally (detailed in B.2). Here, ‘new’ means that the Lyapunov function is not included in the existing literature.
However, it gets the lowest success rate with the 2-D inverted pendulum system and the 6-D 3-bus power system. The challenge comes from optimizing the constants in the Lyapunov function based on given physical constants in the dynamics. Considering the inverted pendulum system, its local Lyapunov function can be represented as $V(x_{1},x_{2})=(1-cos(x_{1}))+\frac{l}{2g}x_{2}^{2}$, where $x_{1}$ is the angular position from the inverted position, $x_{2}$ is the angular velocity, and $g,m,l,b$ are acceleration of gravity, mass of inverted object, the length of string, and the coefficient of friction respectively. Setting $l=5$\mathrm{m}$$, $g=9.81m/s^{2}$, it is challenging for the symbolic regression to find the correct coefficients. In failed cases, our framework often finds the correct structure but incorrect coefficients, e.g. $V=1-cos(x_{1})+0.26x_{2}^{2}$, which is not valid.
In cases where prior dynamics knowledge is available, we incorporate this knowledge into the Lyapunov function design to ease this task. Specifically, for the 4-D rotating wheel pendulum, we use the energy function $E$ as part of the input to the neural Lyapunov function, along with all state variables. As a result, a Lyapunov based on the energy function is revealed by our method. Details of all tested dynamics and considered local regions are provided in Appendix B.
@@ 6 Conclusion
In this paper, we propose CoNSAL, a framework to construct analytical Lyapunov functions for stable nonlinear dynamics. Our approach integrates a neural network learner with a symbolic regression solver, which extracts an analytical formula from the neural Lyapunov function. This symbolic Lyapunov function is further validated with numerical root-finding approaches. When the Lyapunov function is invalid, we efficiently sample counterexamples around the roots to facilitate the training process. The resulting valid Lyapunov function can potentially demonstrate global stability or aid experts in designing controllers. Compared to existing work, CoNSAL provides an analytical Lyapunov function instead of neural ones, which significantly improves the interpretability and generalizability of the final results. Moreover, it bypasses the need for neural network verification, enabling the use of larger neural networks and allowing for scalability to more complex systems. The efficiency of the proposed framework is validated in multiple nonlinear dynamical systems by successfully finding a valid local Lyapunov function for these test cases.
@@
Acknowledgements
This work was supported under grants NSF ECCS-2200692 and Jacobs School Early Career Faculty Development Award 2023. Jie Feng is supported by the UC-National Laboratory In-Residence Graduate Fellowship L24GF7923. Haohan Zou is funded by UC San Diego Halıcıoğlu Data Science Institute (HDSI) scholarship. Yuanyuan Shi is partially supported by the Hellman Fellowship.
Appendix A Code Availability
The code for our proposed algorithm CoNSAL is available at https://github.com/HaohanZou/CoNSAL.
Appendix B Dynamics of the Tested Systems
B.1 Inverted Pendulum
@@ B.2 Van Der Pol Oscillator
B.4 Rotating Wheel Pendulum
@@ B.6 Power System
Appendix C Neural Network Structure and Symbolic Regression Setting